john is two years less than times marty's age. in five years the sum of their ages will be 52. How old is john now?
Answer by ortarico(23)

Jack's age = (x - 2)
Marty's age = (x)
in five years time:
5 + (x - 2) + 5 + (x) = 52
2x + 8 = 52
2x = 44
x = 22
.: jack is 20 yrs old while Marty is 22 yrs old